# Peaceful Anarcho-Accelerationism: Decentralized Full Automation for a Society of Universal Care

### v2

@@ 5 The Liberation Stack: A Technical Architecture
Anarcho-accelerationism is not merely a philosophical position. It proposes a concrete layered architecture for sociotechnical systems that we term the Liberation Stack, illustrated in Figure˜6. Each layer is composed of existing, mature, open-source technologies.
@@
Table˜5 summarizes the frontier energy technologies that underpin the energy layer of the Liberation Stack. These are not speculative proposals; they represent technologies with demonstrated prototypes, substantial investment, and clear pathways to commercial deployment.
@@
The Liberation Stack is designed for compositional sovereignty: each layer can be adopted independently, but the layers reinforce each other. A community with energy sovereignty has more bargaining power. A community with communication sovereignty cannot be silenced. A community with governance tools can make collective decisions about all other layers. Together, the layers form a complete alternative infrastructure that does not depend on either state provision or market availability. Table˜6 summarizes the architecture.
@@ 6 Empirical Evidence: Commons That Work
Founded in 1956 in the Basque Country, Mondragon encompasses approximately 70,000 worker-owners across some 90 cooperatives, a university, 15 research centers, and a social welfare system. With over € 11 billion in annual revenue, it is the seventh-largest business group in Spain (Whyte and Whyte, 1991). Worker-owners vote on major decisions (one person, one vote), elect management, and the pay ratio between highest and lowest earners is capped at 6:1 (compared to 399:1 in large US corporations (Mishel and Kandra, 2021)). Table˜7 presents comparative data on cooperative versus conventional enterprise performance across multiple metrics, demonstrating that the cooperative model is not merely viable but superior on most dimensions.
@@
Table˜8 summarizes the cross-case evidence. These are not marginal experiments. They are large-scale, long-duration, empirically verified systems that embody the principles of anarcho-accelerationism. They demonstrate that a world built on care, cooperation, and commons is not a fantasy; it is a pattern that already exists and can be replicated.
@@
“This is utopian.” The case studies in Section˜6 are not utopian: they are operational systems with decades of track record. Linux is not naive; it runs the internet. Mondragon is not utopian; it employs 70,000 people. Calling a working system “utopian” is not an argument; it is a refusal to engage with evidence. As Ursula K. Le Guin wrote: “We live in capitalism. Its power seems inescapable. So did the divine right of kings” (Le Guin, 2014).
@@ 7 Roadmap: Milestones, Assumptions, and Limitations
A framework without a path of implementation is merely an aspiration. In this section we propose a phased roadmap toward the care-centered society, make explicit the assumptions on which it depends, identify the hard constraints it must respect, and acknowledge its limitations honestly. Table˜9 summarizes the four phases.
@@
Phase I is already underway. The technologies described in the Liberation Stack (Section˜5) are mature, the governance models are proven (Section˜6), and the cost curves for renewable energy and robotic hardware are declining exponentially (Figure˜3). The task of this phase is integration: assembling the existing components into complete, self-sustaining community stacks and demonstrating that they work together. Concretely, this means deploying community solar microgrids with open-source energy management, connecting them to fablabs with robotic manufacturing capability, establishing agroecological food production augmented by autonomous farming robots, linking everything through community-owned mesh networks, governing the whole through open-source democratic platforms, and beginning to experiment with UDR at the local level: residents of pilot communities request goods and services from the productive commons rather than purchasing them on markets.
@@
This roadmap rests on several explicit assumptions. First, we assume that the exponential trajectory of AI and robotic capability continues, as predicted by the universal approximation theorem and DRL convergence results (Section˜2), and as empirically observed over the past two decades (Table˜1). If progress stalls or decelerates significantly, Phases III and IV will be delayed but not invalidated: the Liberation Stack provides value at every stage of automation, not only at full automation. Second, we assume that sufficient renewable energy can be deployed to power full automation. Current projections from IRENA (IRENA, 2023) and the exponential decline in solar costs (Figure˜3) support this assumption, but it requires sustained investment in energy storage, grid infrastructure, and materials recycling. Third, we assume that open-source AI remains competitive with proprietary models. The current ecosystem (LLaMA, Mistral, Stable Diffusion, and their successors) suggests that this is viable, but sustained community investment in training infrastructure is essential. Fourth, we assume that communities can develop effective commons governance at scale. The empirical evidence (Section˜6) supports this for systems up to millions of participants, but planetary-scale federation remains an open institutional design problem.
@@ 8 Conclusion and Future Work
We have proposed the concept of Universal Desired Resources (UDR) as a post-monetary design principle that supersedes Universal Basic Income, and we have shown that UDR achieves the most radical intersectional intervention ever proposed: by eliminating the material basis of oppression, it simultaneously dissolves all axes of structural inequality identified by intersectional theory (Crenshaw, 1989; Collins, 2019), including class, race, gender, disability, sexuality, age, and nationality (Section˜4.1). Drawing on Maslow’s hierarchy of needs (Maslow, 1943), we have demonstrated that the Liberation Stack satisfies basic human needs universally, liberating every person to pursue self-actualization in what we term synthetic liberty: the positive freedom that emerges when commons-based infrastructure provides the material conditions for genuine autonomy, against the merely negative liberty of neoliberalism (Berlin, 1969; Sen, 1999) (Section˜4.2). We have proposed a progressive transition path from UBI to UDR that leverages capitalist technological development while building commons alternatives (Section˜4.3).
We have demonstrated that anarcho-accelerationism respects freedom of faith on principled epistemological and ontological grounds: the interior journey of the soul, as mapped by Saint Teresa of Avila and Saint Augustine (Lumbreras and Garrido-Merchán, 2024), lies categorically beyond the reach of any computational system, and full automation liberates time for the deepest forms of spiritual life (Section˜3). We have established the ethical permissibility of delegating all instrumental labor to machines by demonstrating the independence of phenomenal consciousness from computational intelligence (Garrido-Merchán and Lumbreras, 2022; Garrido-Merchán, 2024; Garrido-Merchán and Blanco, 2022), and we have proposed that moral policy for the commons can be studied and refined through deep reinforcement learning over humanist reward functions (Garrido-Merchán and Lumbreras-Sancho, 2023).
We have proposed the Liberation Stack as a concrete layered architecture for sociotechnical sovereignty, with its energy layer powered by frontier clean energy technologies (perovskite tandem solar, enhanced geothermal, nuclear fusion, iron-air batteries, green hydrogen) within an accelerationist ecological framework that achieves sustainability through technological abundance rather than degrowth (Section˜5). We have presented empirical evidence that commons-based systems already operate at planetary scale (Section˜6), and outlined a phased roadmap with explicit assumptions, constraints, and limitations (Section˜7).

### v3

@@ 6 The Liberation Stack: A Technical Architecture
Anarcho-accelerationism is not merely a philosophical position. It proposes a concrete layered architecture for sociotechnical systems that we term the Liberation Stack, illustrated in Figure˜8. Each layer is composed of existing, mature, open-source technologies.
@@
Table˜6 summarizes the frontier energy technologies that underpin the energy layer of the Liberation Stack. These are not speculative proposals; they represent technologies with demonstrated prototypes, substantial investment, and clear pathways to commercial deployment.
@@
The Liberation Stack is designed for compositional sovereignty: each layer can be adopted independently, but the layers reinforce each other. A community with energy sovereignty has more bargaining power. A community with communication sovereignty cannot be silenced. A community with governance tools can make collective decisions about all other layers. Together, the layers form a complete alternative infrastructure that does not depend on either state provision or market availability. Table˜7 summarizes the architecture.
@@ 0 Energy Energetic sovereignty Solar cooperatives, OpenEMS, microgrids
Injustices Poverty Ecological crisis Techno-feudalism Precarity Intersectional oppression Understanding Reality Econometric evidence Scientific frontiers Historical analysis Commons research Intellectual Commitment Chomsky (Chomsky, 1967): “responsibility of intellectuals” Theory + praxis Social Levers Liberation Stack Open-source commons UBI $\to$ UDR Federated governance Peaceful transition MotivatesInformsActivatesTransforms: eliminates material basis of injustice
[…]
7 Scientific Frontiers for Accelerationist Ecologism
The accelerationist ecological thesis (Section˜6) holds that sustainability is best achieved through technological abundance rather than degrowth. In this section we provide rigorous scientific evidence across four frontiers, namely econometric decoupling, clean energy, matter creation, and biological engineering, demonstrating that the technologies for post-scarcity without ecological destruction are not speculative but empirically grounded.
7.1 Econometric Evidence: Technology Drives Environmental Improvement
The Environmental Kuznets Curve (EKC) hypothesis, first established by Grossman and Krueger (1995) and named by Panayotou (1993), proposes an inverted-U relationship between per capita income and environmental degradation: pollution rises during early industrialization but declines beyond a turning point as societies invest in cleaner technologies and environmental regulation. While Stern (2004) noted important statistical caveats, the most comprehensive systematic review to date, screening 11,500+ papers and analyzing 835 full texts, confirms that relative decoupling of GDP from resource use and emissions is frequent, and absolute decoupling has been achieved in several developed economies for both production-based and consumption-based CO2 emissions (Haberl et al., 2020).
The evidence for decoupling has strengthened dramatically. Le Quéré et al. (2019) analyzed 18 developed economies (representing 28% of global emissions) that achieved sustained emissions reductions between 2005 and 2015, finding that renewable energy deployment is the primary driver. The OECD Green Growth framework (OECD, 2011) documents that green productivity rose faster than traditional Total Factor Productivity across 30 member states. Way et al. (2022) demonstrate that technology cost forecasting based on Wright’s learning curves (Wright, 1936) and experience curves (Lafond et al., 2018) accurately predicts the cost decline trajectory of clean energy technologies, and that a fast transition to renewables would save the global economy approximately \$12 trillion compared to maintaining fossil fuel systems.
Critically, Acemoglu et al. (2012) provide the theoretical foundation for directed technical change: policy interventions that redirect innovation toward clean technologies can permanently alter the trajectory of economic growth away from pollution-intensive paths. Combined with McAfee (2019)’s demonstration of “dematerialization,” the empirical finding that advanced economies are using fewer total resources even as GDP grows, and the World Bank’s framework for inclusive green growth (World Bank, 2012), the econometric evidence strongly supports the accelerationist thesis: the path to ecological sustainability runs through technological acceleration, not degrowth.
7.2 Frontier Clean Energy: Toward Unlimited Clean Power
The convergence of multiple breakthrough energy technologies makes the prospect of abundant, clean, decentralized energy credible for the first time in history.
In nuclear fusion, critical milestones have been crossed. The National Ignition Facility (NIF) achieved target gain $Q>1$ in December 2022, producing 3.15 MJ of fusion energy from 2.05 MJ of laser input, the first laboratory demonstration of fusion ignition (Abu-Shawareb et al., 2024). The SPARC tokamak at MIT, using high-temperature superconducting (REBCO) magnets that achieved a record 20-tesla field, projects $Q\approx 11$ with operations beginning in 2026 (Creely et al., 2020). The Wendelstein 7-X stellarator demonstrated reduced neoclassical energy transport, validating the optimized stellarator concept with record triple product values (Beidler et al., 2021), and Wurzel and Hsu (2022) provide a unified framework for assessing fusion progress across all approaches against the Lawson criterion.
In solar photovoltaics, perovskite-silicon tandem cells have shattered efficiency records. LONGi achieved NREL-certified 33.89% power conversion efficiency in 2024, rising to 34.85% in 2025, the first tandem cell to exceed the single-junction Shockley-Queisser limit (Liu et al., 2024; Green et al., 2024). These cells can be printed on flexible substrates, enabling solar integration into virtually any surface at dramatically reduced manufacturing costs.
Enhanced Geothermal Systems (EGS) have moved from concept to commercial deployment. Fervo Energy demonstrated sustained power generation from engineered reservoirs, with a 400 MW commercial plant under construction (Tester et al., 2006). The US DOE estimates 5,500+ GW of EGS potential in the United States alone, enough to power the entire US economy multiple times over, providing 24/7 carbon-free baseload energy accessible virtually anywhere on Earth.
In multi-day energy storage, iron-air batteries have been revolutionary. Form Energy’s 100+ hour storage technology at projected costs below \$20/kWh uses only iron, water, and air, some of Earth’s most abundant and non-toxic materials (Kwon et al., 2024). Combined with green hydrogen produced at 95%+ efficiency via innovations like H2Pro’s E-TAC electrolysis (Chatterjee et al., 2021), and DOE-backed long-duration storage programs (U.S. Department of Energy, 2024), the intermittency problem of renewables is being solved.
7.3 Creating Matter from Energy: Post-Scarcity Manufacturing
The relationship $E=mc^{2}$ implies that matter and energy are interconvertible. While wholesale matter creation from pure energy remains energetically impractical at current scales, the frontier of advanced manufacturing is converging toward the ability to create virtually any material object from basic feedstocks with minimal waste.
Pair production, the creation of matter-antimatter pairs from photon collisions, has been experimentally demonstrated. Breit and Wheeler (1934) provided the theoretical prediction, Burke et al. (1997) achieved the first laboratory demonstration at SLAC, and the STAR Collaboration at RHIC provided definitive evidence of the linear Breit-Wheeler process with real photons (STAR Collaboration, 2021). While pair production remains a physics demonstration rather than a manufacturing method, it establishes the fundamental principle that matter can be created from energy.
Additive manufacturing is achieving practical matter creation from digital designs. Metal 3D printing achieves buy-to-fly ratios of 1:1 to 3:1, compared to 10:1–30:1 for conventional machining, dramatically reducing material waste (DebRoy et al., 2018). 3D bioprinting fabricates living tissues including skin, bone, vascular grafts, and cartilage (Murphy and Atala, 2014). DNA origami enables atomically precise construction at the nanoscale (Rothemund, 2006), progressing toward Drexler’s vision of molecular manufacturing (Drexler, 1992).
Nuclear transmutation has created entirely new elements that do not exist in nature. Using 48Ca beams on actinide targets, researchers at JINR, GSI, and RIKEN have synthesized elements 107–118 (Oganessian and Utyonkov, 2015), demonstrating that humans can convert one form of matter into another using kinetic energy, genuine transmutation at industrial research scales.
7.4 Biological Engineering: Super Plants and Living Factories
The engineering of photosynthesis represents one of the most promising frontiers for simultaneously addressing energy, carbon capture, and food production.
The RIPE Project (Realizing Increased Photosynthetic Efficiency) has achieved remarkable results. South et al. (2019) engineered a synthetic photorespiratory bypass in tobacco that increased field biomass productivity by over 40%, the largest photosynthetic improvement ever demonstrated in field trials. Kromdijk et al. (2016) accelerated recovery from photoprotection, increasing biomass by 14–20% through a mechanism common to all flowering plants. The C4 Rice Project at IRRI (Ermakova et al., 2020) aims to install C4 photosynthetic machinery into rice, potentially increasing efficiency by 50% in tropical conditions. Salesse-Smith et al. (2018) achieved a 30% increase in RuBisCO content and 15% increase in CO2 assimilation through enzyme overexpression in maize.
[…]
A particularly striking frontier is biophotovoltaics (BPV), which harness photosynthetic organisms to generate electricity directly. McCormick et al. (2015) review the field, documenting how cyanobacteria and algae transfer electrons extracellularly to electrode surfaces. Plant-microbial fuel cells (Strik et al., 2008) generate electricity from living plant root exudates, with the Plant-e technology already commercialized. Engineering CAM photosynthesis into C3 crops (Borland et al., 2014) could reduce water use by up to 80%, enabling food and bioenergy production on arid lands.
7.5 Complementary Sciences: From Quantum Computing to Circular Economy
Several additional scientific frontiers reinforce the accelerationist ecological thesis.
In quantum computing for materials design, the variational quantum eigensolver (Peruzzo et al., 2014) and quantum chemistry algorithms (Cao et al., 2019) enable exponentially faster simulation of molecular behavior, potentially revolutionizing the rational design of catalysts, battery materials, and superconductors without costly trial-and-error synthesis. In nanotechnology, from carbon nanotubes (Iijima, 1991) to graphene (Novoselov et al., 2004), nanomaterials offer revolutionary capabilities for energy storage, water purification, and structural materials at a fraction of the weight and environmental cost of conventional alternatives. Nature’s 3.8 billion years of evolution provide what biomimicry researchers call open-source design blueprints for efficient buildings, drag-reducing surfaces, and ultra-tough materials (Benyus, 1997), democratizing innovation without intellectual property barriers.
The circular economy demonstrates that waste can be eliminated by design. The cradle-to-cradle framework (McDonough and Braungart, 2002) and industrial symbiosis networks like Kalundborg (Jacobsen, 2006) show how inter-firm exchange of water, steam, and by-products can eliminate 635,000 tonnes of CO2 annually, consistent with anarchist principles of emergent, self-organizing order. In direct air capture, Keith et al. (2018) demonstrated industrial DAC at \$94–232/tonne CO2, and Hepburn et al. (2019) review ten pathways for utilizing captured CO2 in fuels, building materials, and chemicals, converting an atmospheric pollutant into feedstock for post-scarcity production. Finally, Drozdov et al. (2019) achieved superconductivity at 250 K ($23\degreeC$) in lanthanum hydride, demonstrating that room-temperature superconductivity is physically achievable, and that practical superconducting power transmission could eliminate the 8–15% of electricity currently lost in grid distribution.
Together, these scientific frontiers demonstrate that the technologies for post-scarcity without ecological destruction are not science fiction. They are active research programs with demonstrated results, declining costs, and clear pathways to deployment. The accelerationist ecological thesis is empirically grounded: we do not need to choose between human flourishing and planetary health. We need to accelerate the right technologies under commons governance.
@@ 8 Empirical Evidence: Commons That Work
Founded in 1956 in the Basque Country, Mondragon encompasses approximately 70,000 worker-owners across some 90 cooperatives, a university, 15 research centers, and a social welfare system. With over € 11 billion in annual revenue, it is the seventh-largest business group in Spain (Whyte and Whyte, 1991). Worker-owners vote on major decisions (one person, one vote), elect management, and the pay ratio between highest and lowest earners is capped at 6:1 (compared to 399:1 in large US corporations (Mishel and Kandra, 2021)). Table˜8 presents comparative data on cooperative versus conventional enterprise performance across multiple metrics, demonstrating that the cooperative model is not merely viable but superior on most dimensions.
@@
Table˜9 summarizes the cross-case evidence. These are not marginal experiments. They are large-scale, long-duration, empirically verified systems that embody the principles of anarcho-accelerationism. They demonstrate that a world built on care, cooperation, and commons is not a fantasy; it is a pattern that already exists and can be replicated.
@@
“This is utopian.” The case studies in Section˜8 are not utopian: they are operational systems with decades of track record. Linux is not naive; it runs the internet. Mondragon is not utopian; it employs 70,000 people. Calling a working system “utopian” is not an argument; it is a refusal to engage with evidence. As Ursula K. Le Guin wrote: “We live in capitalism. Its power seems inescapable. So did the divine right of kings” (Le Guin, 2014).
@@ 9 Roadmap: Milestones, Assumptions, and Limitations
A framework without a path of implementation is merely an aspiration. In this section we propose a phased roadmap toward the care-centered society, make explicit the assumptions on which it depends, identify the hard constraints it must respect, and acknowledge its limitations honestly. Table˜10 summarizes the four phases.
@@
Phase I is already underway. The technologies described in the Liberation Stack (Section˜6) are mature, the governance models are proven (Section˜8), and the cost curves for renewable energy and robotic hardware are declining exponentially (Figure˜3). The task of this phase is integration: assembling the existing components into complete, self-sustaining community stacks and demonstrating that they work together. Concretely, this means deploying community solar microgrids with open-source energy management, connecting them to fablabs with robotic manufacturing capability, establishing agroecological food production augmented by autonomous farming robots, linking everything through community-owned mesh networks, governing the whole through open-source democratic platforms, and beginning to experiment with UDR at the local level: residents of pilot communities request goods and services from the productive commons rather than purchasing them on markets.
@@
This roadmap rests on several explicit assumptions. First, we assume that the exponential trajectory of AI and robotic capability continues, as predicted by the universal approximation theorem and DRL convergence results (Section˜2), and as empirically observed over the past two decades (Table˜1). If progress stalls or decelerates significantly, Phases III and IV will be delayed but not invalidated: the Liberation Stack provides value at every stage of automation, not only at full automation. Second, we assume that sufficient renewable energy can be deployed to power full automation. Current projections from IRENA (IRENA, 2023) and the exponential decline in solar costs (Figure˜3) support this assumption, but it requires sustained investment in energy storage, grid infrastructure, and materials recycling. Third, we assume that open-source AI remains competitive with proprietary models. The current ecosystem (LLaMA, Mistral, Stable Diffusion, and their successors) suggests that this is viable, but sustained community investment in training infrastructure is essential. Fourth, we assume that communities can develop effective commons governance at scale. The empirical evidence (Section˜8) supports this for systems up to millions of participants, but planetary-scale federation remains an open institutional design problem.
@@ 10 Conclusion and Future Work
We have proposed the concept of Universal Desired Resources (UDR) as a post-monetary design principle that supersedes Universal Basic Income, and we have shown that UDR achieves the most radical intersectional intervention ever proposed: by eliminating the material basis of oppression, it simultaneously dissolves all axes of structural inequality identified by intersectional theory (Crenshaw, 1989; Collins, 2019), including class, race, gender, disability, sexuality, age, and nationality (Section˜5.1). Drawing on Maslow’s hierarchy of needs (Maslow, 1943), we have demonstrated that the Liberation Stack satisfies basic human needs universally, liberating every person to pursue self-actualization in what we term synthetic liberty: the positive freedom that emerges when commons-based infrastructure provides the material conditions for genuine autonomy, against the merely negative liberty of neoliberalism (Berlin, 1969; Sen, 1999) (Section˜5.2). We have proposed a progressive transition path from UBI to UDR that leverages capitalist technological development while building commons alternatives (Section˜5.3).
We have demonstrated that anarcho-accelerationism respects freedom of faith on principled epistemological and ontological grounds: the interior journey of the soul, as mapped by Saint Teresa of Avila and Saint Augustine (Lumbreras and Garrido-Merchán, 2024), lies categorically beyond the reach of any computational system, and full automation liberates time for the deepest forms of spiritual life (Section˜4). We have established the ethical permissibility of delegating all instrumental labor to machines by demonstrating the independence of phenomenal consciousness from computational intelligence (Garrido-Merchán and Lumbreras, 2022; Garrido-Merchán, 2024; Garrido-Merchán and Blanco, 2022), and we have proposed that moral policy for the commons can be studied and refined through deep reinforcement learning over humanist reward functions (Garrido-Merchán and Lumbreras-Sancho, 2023).
We have proposed the Liberation Stack as a concrete layered architecture for sociotechnical sovereignty, with its energy layer powered by frontier clean energy technologies (perovskite tandem solar, enhanced geothermal, nuclear fusion, iron-air batteries, green hydrogen) within an accelerationist ecological framework that achieves sustainability through technological abundance rather than degrowth (Section˜6). We have presented empirical evidence that commons-based systems already operate at planetary scale (Section˜8), and outlined a phased roadmap with explicit assumptions, constraints, and limitations (Section˜9).